OpenStudy (anonymous):

If v1 = (2,4) and v2 = (-1,5), then v1*v2 is..

OpenStudy (anonymous):

...(-2,20)?

jimthompson5910 (jim_thompson5910):

assuming that's the dot product v1 dot v2 = 2*(-1) + 4*5 v1 dot v2 = -2 + 20 v1 dot v2 = 18 Note: the result of a dot product is ALWAYS a scalar (and not a vector)

OpenStudy (anonymous):

Thank you... so much!

jimthompson5910 (jim_thompson5910):

yw